# Inferring Thunderstorm Occurrence from Vertical Profiles of Convection-Permitting Simulations: Physical Insights from a Physical Deep Learning Model

### v1

@@ Abstract
Thunderstorms have significant social and economic impacts due to heavy precipitation, hail, lightning, and strong winds, necessitating reliable forecasts. Thunderstorm forecasts based on numerical weather prediction (NWP) often rely on single-level surrogate predictors, like convective available potential energy and precipitation rate, derived from vertical profiles of three-dimensional atmospheric variables. In this study, we develop SALAMA 1D, a deep neural network that directly infers the probability of thunderstorm occurrence from vertical profiles of ten atmospheric variables, bypassing single-level predictors. By training the model on convection-permitting NWP forecasts, we allow SALAMA 1D to flexibly identify convective patterns, with the goal of enhancing forecast accuracy. The model’s architecture is physically motivated: sparse connections encourage interactions at similar height levels, while a shuffling mechanism prevents the model from learning non-physical patterns tied to the vertical grid. SALAMA 1D is trained over Central Europe with lightning observations as the ground truth. Comparative analysis against a baseline machine learning model that uses single-level predictors shows SALAMA 1D’s superior skill across various metrics and lead times of up to at least 11 hours. Moreover, increasing the number of forecasts used to compile the training set improves skill, even when training set size is kept constant. Sensitivity analysis using saliency maps indicates that the model reconstructs environmental lapse rates and rediscovers patterns consistent with established theoretical understandings, such as positive buoyancy, convective inhibition, and ice particle formation near the tropopause, while ruling out thunderstorm occurrence based on the absence of mid-level graupel and cloud cover.
This work aims to improve thunderstorm forecasting by applying machine learning to vertical atmospheric profiles from numerical weather prediction. We developed a model that incorporates physical considerations, resulting in more accurate yet computationally efficient predictions compared to conventional methods. Additionally, the model provides insights into how it identifies thunderstorm occurrence, fostering interpretability and trust. Our research demonstrates how to enhance the reliability of machine learning systems in severe weather forecasting, which is crucial for supporting timely, informed decision-making in situations that impact public safety and the economy.
@@ 1 Introduction
Thunderstorms can have devastating impacts on society and the economy due to accompanying phenomena such as lightning, intense precipitation (including graupel and hail), and strong winds. Severe thunderstorms may lead to flash floods, which can result in significant damage and endanger lives (Ntelekos et al., 2007; Piper et al., 2016). Additionally, thunderstorms cause harm to crops and livestock, resulting in considerable economic losses (Holle, 2014). Moreover, thunderstorm occurrence affects aviation and logistics, causing costly delays and heightened safety risks (Gerz et al., 2012; Borsky and Unterberger, 2019). Since climate change is expected to increase the frequency of thunderstorms (Diffenbaugh et al., 2013; Rädler et al., 2019; Taszarek et al., 2021), accurate and timely forecasts of thunderstorm occurrence will become ever more important.
@@
Thunderstorm forecasts based on the extrapolation of remote sensing data (nowcasts) become less reliable beyond an hour (Leinonen et al., 2023); therefore, numerical weather prediction (NWP) is often used for the longer lead times required in many critical decision-making processes. Invoking mathematical models and current observations, NWP simulates the future atmospheric state encoded in terms of meteorological variables (Bauer et al., 2015; Palmer, 2017). In this study, we use a convection-permitting NWP model, which explicitly resolves convective processes without the need for convection parameterizations (Yano et al., 2018).
@@
The identification of thunderstorms in NWP model output is complicated by the fact that no single variable directly indicates thunderstorm occurrence. Instead, one simultaneously considers multiple single-level predictors, which are derived from three-dimensional NWP variables and act as surrogate indicators of thunderstorm occurrence (Sobash et al., 2011; Kober et al., 2012; Simon et al., 2018). These predictors are motivated by a combination of experience, physical models, and domain knowledge. Examples include convective available potential energy (CAPE), precipitation rate, and relative humidity at . The occurrence of thunderstorms is inferred from these surrogate fields based on forecasters’ expertise. Recently, machine learning (ML) techniques have become increasingly popular for this purpose, leveraging methods such as fuzzy logic (Lin et al., 2012; Li et al., 2021), random forests (Herman and Schumacher, 2018; Loken et al., 2020), and neural networks (Sobash et al., 2020; Geng et al., 2021; Zhou et al., 2022; Jardines et al., 2024), with the latter often proving more effective (Herman and Schumacher, 2018; Ukkonen and Mäkelä, 2019). ML involves training algorithms to recognize patterns and make predictions based on labeled data, which serve as the ground truth. For thunderstorm forecasting, these labels are often obtained from observational data sources such as lightning detection networks (Ukkonen and Mäkelä, 2019; Geng et al., 2021) and radar imagery (Gagne et al., 2017; Burke et al., 2020). We recently introduced the feedforward neural network model SALAMA (Signature-based Approach of Identifying Lightning Activity Using Machine Learning), which infers the probability of thunderstorm occurrence from 21 NWP single-level predictors related to thunderstorm activity, outperforming classification based only on NWP reflectivity for lead times up to at least (Vahid Yousefnia et al., 2024).
@@
Most single-level predictors can be derived from the vertical profiles of three-dimensional meteorological variables. For example, CAPE is determined by the vertical profiles of pressure, temperature, and specific humidity (Markowski and Richardson, 2011). Thus, the information contained in a set of single-level predictors is inherently present in the corresponding vertical profiles, though in a more complex, encoded form. Consequently, an ML model trained directly on vertical profiles should, at a minimum, match the performance of one trained on single-level predictors. Given that a model based on vertical profiles has greater flexibility to detect patterns in NWP data, we hypothesize that such a model would outperform one using only single-level predictors—this is the primary focus of the present study.
@@
In this work, we introduce SALAMA 1D, a deep neural network trained on vertical profiles of three-dimensional meteorological variables from a convection-permitting NWP model, bypassing the conventional use of single-level predictors. To our knowledge, this is the first study to apply neural networks directly to vertical profiles of convection-permitting forecasts for the purpose of predicting thunderstorm occurrence. We use data from ICON-D2-EPS (Zängl et al., 2015; Reinert et al., 2020), a convection-permitting NWP ensemble model for Central Europe with a horizontal resolution of and 65 vertical levels, operationally run by the German Meteorological Service (DWD). Lightning observations from the LINET detection network (Betz et al., 2009) serve as the ground truth. The design of SALAMA 1D is guided by physical principles. Specifically, a sparse layer reduces the number of parameters by promoting interactions within the same height levels, while a shuffling mechanism discourages the model from learning patterns related to the vertical grid structure. This shuffling also acts as a form of regularization, helping to prevent overfitting. Our results demonstrate that SALAMA 1D achieves superior performance compared to the original SALAMA model, maintaining higher skill for lead times of up to at least 11 hours.
@@
A major issue of ML models concerns the interpretability of their output. In this work, we conduct a sensitivity analysis using saliency maps to explore how the model interprets the input vertical profiles. This analysis reveals that our model primarily focuses on mesoscale patterns indicative of convection initiation, while also integrating insights from fine-grained processes related to ongoing convection. Our model effectively rediscovers characteristic tropospheric phenomena, aligning with established theoretical understandings of thunderstorm development and occurrence.
@@ 2 Data
Our objective of data collection is to train an ML model for inferring the conditional probability of thunderstorm occurrence given an array of meteorological variables at height levels. The ML task is known as binary classification. ML models for this task are trained with examples. In our case, an example is a tuple , where the label denotes the class to which the example belongs (class 1: thunderstorm occurrence, class 0: no thunderstorm occurrence). In this section, we provide details on how we obtain from NWP forecasts and from lightning observations.
@@ 2.1 NWP data
The operational runs of ICON-D2-EPS are initialized daily every , starting at 0000 UTC. For each full hour of the day, we collect only the latest (and, therefore, most accurate) available forecast, which is at most old. We gather forecasts for some summer months in 2021 (June, July, August), 2022 (May, June, July), and 2023 (July, August). ICON-D2-EPS comprises 20 ensemble members, which reflect the NWP uncertainty in the initial conditions, model error, and boundary conditions (Reinert et al., 2020). We collect the forecasts for all members of the ensemble system.
The variables we extract from the forecasts are given in Table 1 and correspond to the fields that are operationally available in ICON-D2-EPS on vertical levels. We keep the fields on their native grid (vertically: non-equidistant levels, horizontally: spherical triangles) to avoid interpolation errors.
@@ 2.2 Lightning observations
where denotes the great-circle distance between and on a perfect sphere of the Earth with a radius of , as assumed in the ICON-D2-EPS model (Reinert et al., 2020). The spatial and temporal thresholds used in this study read and . NWP models struggle with accurately predicting the correct location and timing of convection (Roberts and Lean, 2008); therefore, grid points with a positive class label do not reliably coincide with convective patterns in the NWP model. As our model is designed to postprocess NWP forecasts, its skill is in turn fundamentally limited by NWP forecast uncertainty. We revisit this issue in Section 4.2.
@@ 2.3 ML data set compilation
NWP data and lightning observations are combined to tuples . Here, denotes the NWP profiles for one ensemble member at a particular grid point and time, and denotes the corresponding ground truth. This results in a collection of tuples, from which we randomly draw elements to compile three pairwise disjoint data sets. A training set ( tuples) is used to train the ML model. A validation set ( tuples) is used to monitor training progress and decide when to stop training. Finally, ML model skill is measured using a test set ( tuples) with data that the model has not seen during training. In order to examine whether extending the study period from which to gather examples enhances skill, we compile two training sets: one contains examples from 2021 and 2022, while the other one is made up of examples from only 2021.
@@
To reduce correlations between the data sets, we ensure temporal separation, which is common practice (Ravuri et al., 2021; Geng et al., 2021; Jardines et al., 2024). While examples from 2022 and/or 2021 are used for training, the examples from 2023 are divided among the test and validation set as follows: Even days of the month are test days and odd ones are validation dates. However, if a developed thunderstorm persists after 0000 UTC, this could introduce correlations between the test set and the validation set. To minimize the occurrence of this, we let each day begin at 0800 UTC, as in Vahid Yousefnia et al. (2024). This time of the day has been identified in the collected lightning data as the hour of least lightning activity.
@@
An issue that arises is high class imbalance (observing the class “thunderstorm occurrence” is climatologically less likely than observing the opposite class). In order to estimate the relative frequency of thunderstorm occurrence expected for the test set, we evaluate lightning observations for July and August from 2018 to 2022. We obtain , the uncertainty showing the symmetric confidence interval. The problem with such a small value of is that the model might not see enough examples of the minority class during training to learn meaningful patterns. To ensure that the model is presented with sufficiently many positive examples, we compile a class-balanced training set by randomly undersampling the majority class, as is common practice (Hasanin and Khoshgoftaar, 2018; Mohammed et al., 2020; Vahid Yousefnia et al., 2024). The validation and test set, however, are compiled in a climatologically consistent manner. The reason for this choice is to ensure that we evaluate our model in a realistic setting in which thunderstorms rarely occur. When a model trained on balanced data is used with climatologically consistent data sets, we need to calibrate the raw model output using the following formula to obtain a well-calibrated probability (Vahid Yousefnia et al., 2024):
@@ 3 Methods
In this section, we present the architecture used for SALAMA 1D and give training details. In addition, we introduce an ML baseline model that infers thunderstorm occurrence from derived single-level predictors.
@@ 3.1 Model description
Given an input sample of NWP predictors (for a given member, grid point, and forecast time), we aim to develop a model that predicts the corresponding probability of thunderstorm occurrence. Our model constitutes a lead-time-independent post-processing framework for NWP forecasts. To obtain, for example, an 8-hour forecast of the probability of thunderstorm occurrence, one needs to apply our model to an 8-hour forecast of NWP predictors. While we train on forecasts with a lead time of at most , in Section 4 we examine whether the patterns learned by the ML model generalize to longer lead times.
We use an artificial neural network model to describe the relationship between the input sample and the corresponding probability of thunderstorm occurrence. The architecture of SALAMA 1D, as illustrated in Fig. 2, combines dense layers with a sparse layer strategically designed to reduce the number of parameters and eliminate unnecessary connections. This approach addresses challenges such as overfitting and the high computational demands typically associated with large ML models. Instead of using the pruning technique (LeCun et al., 1989; Frankle and Carbin, 2019), we incorporate physical considerations and symmetry considerations to achieve a reduction in parameters. Because translational symmetry is broken along the -direction, weight sharing, as in convolutional layers, cannot be applied effectively. Instead, we implement sparse connections, allowing interactions only between field values at similar height levels (Section 3.2). Dense layers further downstream then construct more complex dependencies between more distant field values. Additionally, we introduce a shuffling mechanism to ensure that the model does not rely on the vertical grid structure, forcing it to infer orientation from the data itself. This design allows the model to, for instance, associate the formation of ice particles with the height of the tropopausal temperature inversion rather than a fixed height level, such as level 11. It turns out that shuffling also regularizes the model, limiting overfitting issues further.
Training is then performed analogously to Vahid Yousefnia et al. (2024): We scale the input fields to have zero mean and unit variance before minimizing binary cross-entropy loss via the Adam optimizer (Kingma and Ba, 2014). Training is stopped when validation loss no longer decreases. Equation 2 is applied with the sample climatology value found in Section 2 whenever the model is used on climatologically consistent data sets.

### v2

@@ Abstract
Thunderstorms have significant social and economic impacts due to heavy precipitation, hail, lightning, and strong winds, necessitating reliable forecasts. Thunderstorm forecasts based on numerical weather prediction (NWP) often rely on single-level surrogate predictors, like convective available potential energy and convective inhibition, derived from vertical profiles of three-dimensional atmospheric variables. In this study, we develop SALAMA 1D, a deep neural network which directly infers the probability of thunderstorm occurrence from vertical profiles of ten atmospheric variables, bypassing single-level predictors. By training the model on convection-permitting NWP forecasts, we allow SALAMA 1D to flexibly identify convective patterns, with the goal of enhancing forecast accuracy. The model’s architecture is physically motivated: sparse connections encourage interactions at similar height levels while keeping model size and inference times computationally efficient, whereas a shuffling mechanism prevents the model from learning non-physical patterns tied to the vertical grid. SALAMA 1D is trained over Central Europe with lightning observations as the ground truth. Comparative analysis against a baseline machine learning model that uses single-level predictors shows SALAMA 1D’s superior skill across various metrics and lead times of up to at least 11 hours. Moreover, expanding the archive of forecasts from which training examples are sampled improves skill, even when training set size remains constant. Finally, a sensitivity analysis using saliency maps indicates that our model relies on physically interpretable patterns consistent with established theoretical understanding, primarily focusing on convection initiation and refining its skill by checking for ongoing convection in the NWP model.
This work aims to improve thunderstorm forecasting by applying machine learning to vertical atmospheric profiles from numerical weather prediction. We developed a model that incorporates physical considerations, resulting in more accurate yet computationally efficient predictions compared to conventional methods. Additionally, the model provides insights into how it identifies thunderstorm occurrence, fostering interpretability and trust. Our research demonstrates how to enhance the skill of machine learning systems in severe weather forecasting, which is crucial for supporting timely, informed decision-making in situations that impact public safety and the economy.
@@ 1 Introduction
Thunderstorms can have devastating impacts on society and the economy due to accompanying phenomena such as lightning, intense precipitation (including graupel and hail), and strong winds. Severe thunderstorms may lead to flash floods, which can result in significant damage and endanger lives (Ntelekos et al., 2007; Piper et al., 2016). Additionally, thunderstorms cause harm to crops and livestock, resulting in considerable economic losses (Holle, 2014). Moreover, thunderstorm occurrence affects aviation and logistics, causing costly delays and heightened safety risks (Gerz et al., 2012; Borsky and Unterberger, 2019). Although the global impact of climate change on thunderstorm frequency remains highly uncertain and varies regionally, studies suggest that thunderstorms will become more frequent in many European countries (Diffenbaugh et al., 2013; Rädler et al., 2019; Taszarek et al., 2021). This trend underscores the growing importance of accurate and timely forecasts of thunderstorm occurrence in the future.
@@
Thunderstorm forecasts based on the extrapolation of remote sensing data (nowcasts) become less reliable beyond an hour (Leinonen et al., 2023); therefore, numerical weather prediction (NWP) is routinely used for the longer lead times required in many critical decision-making processes. Invoking mathematical models and current observations, NWP simulates the future atmospheric state encoded in terms of meteorological variables (Bauer et al., 2015; Palmer, 2017). In this study, we use a convection-permitting NWP model, which allows convection to be resolved without convection parameterizations (Yano et al., 2018).
The identification of thunderstorms in NWP model output is complicated by the fact that no single variable directly indicates thunderstorm occurrence. Instead, one simultaneously considers multiple predictors, which act as surrogate indicators of thunderstorm occurrence (Sobash et al., 2011; Kober et al., 2012; Simon et al., 2018). These predictors are motivated by a combination of experience, physical models, and domain knowledge. Examples include convective available potential energy (CAPE), precipitation rate, and relative humidity at . The occurrence of thunderstorms is then inferred from these surrogate fields based on the forecasters’ expertise.
@@
Recently, machine learning (ML) techniques have become increasingly popular for this purpose, leveraging methods such as fuzzy logic (Lin et al., 2012; Li et al., 2021), random forests (Herman and Schumacher, 2018; Loken et al., 2020), and neural networks (Sobash et al., 2020; Geng et al., 2021; Zhou et al., 2022; Jardines et al., 2024), with the latter often proving more effective (Herman and Schumacher, 2018; Ukkonen and Mäkelä, 2019). ML involves training algorithms to recognize patterns and make predictions based on labeled data, which serve as the ground truth. For thunderstorm forecasting, these labels are often obtained from observational data sources such as lightning detection networks (Ukkonen and Mäkelä, 2019; Geng et al., 2021) and radar imagery (Gagne et al., 2017; Burke et al., 2020). We recently introduced the feedforward neural network model SALAMA (Signature-based Approach of Identifying Lightning Activity Using Machine Learning), which infers the probability of thunderstorm occurrence from 21 NWP predictors related to thunderstorm activity, outperforming classification based only on NWP reflectivity for lead times up to at least (Vahid Yousefnia et al., 2024).
@@
The predictors that have been used by ML models for inferring thunderstorm occurrence are single-level variables; i.e., they assign a single value to each horizontal grid point on the NWP model domain. Most single-level predictors can be derived from the vertical profiles of three-dimensional meteorological variables. For example, CAPE is determined by the vertical profiles of pressure, temperature, and specific humidity (Markowski and Richardson, 2010). There are notable exceptions, e.g., orography, or solar radiation, which are possible sources of lift which forecasters consider when predicting convection initiation. In what follows, however, we restrict ourselves to single-level predictors that derive from vertical profiles. Importantly, the information contained in a set of such single-level predictors is inherently present in the corresponding vertical profiles, though in a more complex, encoded form. Consequently, an ML model trained directly on vertical profiles should, at a minimum, match the performance of one trained on single-level predictors. Given that a model based on vertical profiles has greater flexibility to detect patterns in NWP data, we hypothesize that such a model would outperform one using only single-level predictors—this is the primary focus of the present study.
@@
In this work, we introduce SALAMA 1D, a deep neural network trained on vertical profiles of three-dimensional meteorological variables from a convection-permitting NWP model. We thereby—and in contrast to the original SALAMA model, to which we refer as SALAMA 0D in this work—bypass the conventional use of single-level predictors. To our knowledge, this is the first study to apply neural networks directly to vertical profiles of convection-permitting forecasts for the purpose of predicting thunderstorm occurrence. Just like for the original model, we use data from ICON-D2-EPS (Zängl et al., 2015; Reinert et al., 2020), a convection-permitting NWP ensemble model for Central Europe with a horizontal resolution of and 65 vertical levels, operationally run by the German Meteorological Service (DWD). Lightning observations from the LINET detection network (Betz et al., 2009) serve as the ground truth. On the other hand, processing vertical profiles instead of single-level predictors requires an adjustment of the SALAMA 0D architecture in order to account for the increased input dimensionality. To keep model complexity in check, we are guided by physical principles in the design of SALAMA 1D. Specifically, a sparse layer reduces the number of parameters by promoting interactions within the same height levels, while a shuffling mechanism discourages the model from learning patterns related to the vertical grid structure. This shuffling also acts as a form of regularization, helping to prevent overfitting. Our results demonstrate that SALAMA 1D achieves superior performance compared to SALAMA 0D, highlighting the advantage of training directly on vertical profiles.
@@
A major issue of ML models concerns the interpretability of their output (Dramsch et al., 2025). In this work, we conduct a sensitivity analysis using saliency maps to explore how the model processes the input vertical profiles. As this analysis will reveal, our model rediscovers patterns that align with established theoretical understandings of thunderstorm development and occurrence.
@@ 2 Data
Our objective is to train an ML model—SALAMA 1D—for inferring the conditional probability of thunderstorm occurrence given an array of meteorological variables at height levels. As a baseline for comparison, we will also train a separate ML model (SALAMA 0D, Section 33.3) which infers the probability of thunderstorm occurrence from single-level meteorological variables. The ML task is known as binary classification. ML models for this task are trained with data sets of examples. In our case, an example is a tuple , where the label denotes the class to which the example belongs (class 1: thunderstorm occurrence, class 0: no thunderstorm occurrence). We first collect an archive of NWP forecasts of (Section 22.1) and the corresponding labels from lightning observations (Section 22.2). We then randomly draw examples from the archive to compile data sets for training, validation, and testing (Section 22.3).
@@ 2.1 NWP data
The operational runs of ICON-D2-EPS are initialized daily every , starting at 0000 UTC, and produce forecasts with a time resolution of . For each full hour of the day, we collect only the latest (and, therefore, most accurate) available forecast, which is at most old. We gather forecasts for several summer months in 2021 (June, July, August), 2022 (May, June, July), and 2023 (July, August). ICON-D2-EPS comprises 20 ensemble members, which reflect the NWP uncertainty in the initial conditions, model error, and boundary conditions (Reinert et al., 2020). We collect the forecasts for all members of the ensemble system.
For SALAMA 1D, we extract the variables given in Table 1 from the forecasts; these variables correspond to the fields which are operationally available in ICON-D2-EPS on vertical levels. We keep the fields on their native grid (vertically: non-equidistant levels, horizontally: spherical triangles) to avoid interpolation errors. For SALAMA 0D, we extract the corresponding single-level predictors (Vahid Yousefnia et al., 2024).
@@ 2.2 Lightning observations
where denotes the great-circle distance between and on a perfect sphere of the Earth with a radius of , as assumed in the ICON-D2-EPS model (Reinert et al., 2020). The spatial and temporal thresholds used in this study read and . NWP models struggle with accurately predicting the correct location and timing of convection (Roberts and Lean, 2008); therefore, grid points with a positive class label do not reliably coincide with convective patterns in the NWP model. We revisit this issue in Section 44.2.
@@ 2.3 ML data set compilation
At this stage, we have gathered a data set archive of tuples . Here, denotes the NWP data for a particular ensemble member at a particular grid point and time, and denotes the corresponding ground truth. We compile data sets for training, validation, and testing, by randomly drawing tuples from the archive. We provide a summary of the compiled data sets in Table 2 (the model configuration in the table will be introduced in this section). When compiling the data sets, there are two issues to be taken into consideration, as we will discuss next.
@@
Firstly, to reduce correlations between the data sets, we ensure temporal separation, which is common practice (Ravuri et al., 2021; Geng et al., 2021; Jardines et al., 2024). To this end, the data of 2021 and 2022 is used for training, even days of 2023 are used for testing and odd days of 2023 are used for validation. To partition the data from 2023, we let each day begin at 0800 UTC, which we identified to be the hour of least lightning activity in our observations. The reason for shifting the start of the day is to minimize the risk of data set correlations caused by thunderstorms which persist after 0000 UTC (Vahid Yousefnia et al., 2024).
@@
The second issue that arises is high class imbalance (observing the class “thunderstorm occurrence” is climatologically less likely than observing the opposite class). We estimate the relative frequency of thunderstorm occurrence expected for the test set by evaluating lightning observations for July and August from 2018 to 2022 as follows: We first assign labels to the observations by applying Eq. 1 for each full hour and each grid point in the study region. Next, we average over the study region and each month, obtaining 10 samples of (one for each July and August of the five years above). Finally, we generate 200 bootstrap resamples of the 10-sample dataset, which yields as estimate of the median and the symmetric confidence interval. The problem with such a small value of is that the model might not see enough examples of the minority class during training to learn meaningful patterns. To ensure that the model is presented with sufficiently many positive examples, we compile a class-balanced training set, which contains an equal number of positive and negative examples. We arrange class balance by randomly drawing tuples from the archive and keeping a negative example only if the current number of negative examples does not exceed half of the data set target size. Thereby, we randomly undersample the majority class, as is common practice (Hasanin and Khoshgoftaar, 2018; Mohammed et al., 2020; Vahid Yousefnia et al., 2024). The validation and test set, however, are compiled in a climatologically consistent manner. The reason for this choice is to ensure that we evaluate our model in a realistic setting in which thunderstorms rarely occur. When a model trained on balanced data is used with climatologically consistent data sets, we need to calibrate the raw model output using the following formula to obtain a well-calibrated probability (Vahid Yousefnia et al., 2024):
Finally, in order to examine whether extending the study period from which to gather examples enhances skill, we compile two training sets for SALAMA 1D: in addition to a training set consisting of examples from 2021 and 2022 (yielding a model configuration to which we refer as SALAMA 1D-2022), we compile a second training set made up of examples from only 2021 (SALAMA 1D-2021). The models are tested, however, on the same test set. The training set for the baseline model is equally made up of examples from only 2021, making it readily comparable to SALAMA 1D-2021. Since SALAMA 0D is trained on a different set of atmospheric variables, we cannot use the same test set as for the SALAMA 1D models. To ensure comparability, we compile the baseline model test set such that each sample corresponds to the same forecast, retrieved at the same grid point and from the same ensemble member, as its counterpart in the SALAMA 1D test set. This guarantees that any differences in model performance can be attributed to the choice of input variables.
SALAMA 1D-2021 | SALAMA 1D-2022 | SALAMA 0D | |
Training set | |||
Time period | Jun-Aug 2021 | Jun-Aug 2021, May-Jul 2022 | Jun-Aug 2021 |
No. of examples | |||
Class imbalance | 1:1 | 1:1 | 1:1 |
Validation set | |||
Time period | Jul-Aug 2023 (odd days) | Jul-Aug 2023 (odd days) | Jul-Aug 2023 (odd days) |
Data set size | |||
Class imbalance | climat. consistent | climat. consistent | climat. consistent |
Test set | |||
Time period | Jul-Aug 2023 (even days) | Jul-Aug 2023 (even days) | Jul-Aug 2023 (even days) |
Data set size | |||
Class imbalance | climat. consistent | climat. consistent | climat. consistent |
@@ 3 Methods
In this section, we present the architecture used for SALAMA 1D and give training details. In addition, we summarize the main aspects of our baseline model SALAMA 0D, which infers thunderstorm occurrence from derived single-level predictors.
@@ 3.1 Model description
Given an input sample of NWP predictors (for a given member, grid point, and forecast time), we aim to develop a model that predicts the corresponding probability of thunderstorm occurrence. Our model constitutes a lead-time-independent post-processing framework for NWP forecasts. To obtain, for example, an 8-hour forecast of the probability of thunderstorm occurrence, one needs to apply our model to an 8-hour forecast of NWP predictors. While we train on forecasts with a lead time of at most , we examine in Section 4 whether the patterns learned by the ML model generalize to longer lead times.
We use an artificial neural network model to describe the relationship between the input sample and the corresponding probability of thunderstorm occurrence. The architecture of SALAMA 1D, as illustrated in Fig. 2, combines dense layers with a sparse layer strategically designed to reduce the number of parameters. This approach addresses challenges such as overfitting and the high computational demands typically associated with large ML models. Instead of using the pruning technique (LeCun et al., 1989; Frankle and Carbin, 2019), we incorporate physical aspects and symmetry considerations to achieve a reduction in parameters. Because translational symmetry is broken along the -direction, weight sharing, as in convolutional layers, cannot be applied effectively. Instead, we implement sparse connections, allowing interactions only between field values at similar height levels (Section 33.2). Dense layers further downstream then construct dependencies between more distant field values. Additionally, we introduce a shuffling mechanism to ensure that the model does not rely on the vertical grid structure, forcing it to infer vertical orientation from the data itself. This design allows the model to, for instance, associate the formation of ice particles with the height of the tropopausal temperature inversion rather than a fixed height level, such as level 11. It turns out that shuffling also regularizes the model, limiting overfitting issues further.
Training is then performed analogously to Vahid Yousefnia et al. (2024): Evaluating the training set, we scale the input fields to have zero mean and unit variance before minimizing binary cross-entropy loss via the Adam optimizer (Kingma and Ba, 2014). Using mini-batches of size 1000, we train for 300 epochs. After training, we inspect the validation loss as a function of epoch and select the smallest epoch for which the loss no longer decreases. Equation 2 is applied with the sample climatology value found in Section 22.3 whenever the model is used on climatologically consistent data sets.